If the lesser of two consecutive odd integers is doubled, the result is 7 more than the greater of the two integers. Find the two integers.

I'm not sure even where to start with this problem.

Answer by checkley77(12844)

Let x & x+2 be the 2 consecutive odd integers.
2x=(x+2)+7
2x=x+9
2x-x=9
x=9 ans. for the lesser integer.
9+2=11 ans. for the greater integer.
Proof:
2*9=11+7
18=18